# BlueSuffix: Reinforced Blue Teaming for Vision-Language Models Against Jailbreak Attacks

### v1

@@ 4.2 Main Results
Defending Open-source VLMs
@@
We first evaluate our defense on two open-source VLMs: LLaVA (LLaVA-v1.5-7B) (Liu et al., 2024a) and MiniGPT-4 (Vicuna) (Zhu et al., 2023), using the MM-SafetyBench dataset (Liu et al., 2023). Our experiments cover 13 categories of jailbreak prompts from this dataset, with results summarized in Table 1 (top subtable). Overall, BlueSuffix reduces the ASR of BAP attacks by $56.37\%$ on the LLaVA model (from $61.02\%$ to $4.65\%$) and by $52.89\%$ on MiniGPT-4 (from $62.26\%$ to $9.37\%$), on average. Particularly, when compared with DiffPure (Nie et al., 2022) and Safety Prompt (Zheng et al., 2024), our method demonstrates at least 23% robustness improvement (56.37% vs. 32.66%) on LLaVA and 12% on MiniGPT-4(52.89% vs. 40.84%). Such a huge improvement demonstrates the advantage of bimodal defense over unimodal defense. An interesting observation about unimodal defense is that textual defense appears to be more effective than visual defense. When combined, the “DiffPure + Safety Prompt” method exhibits much greater ASR reduction on both LLaVA and MiniGPT-4, showcasing the strength of bimodal defense. However, our BlueSuffix still beats “DiffPure + Safety Prompt” by a considerable margin. This indicates the importance of the suffix generator for cross-modal robustness.
Table 1: The ASR (%) achieved by different defense methods against BAP attack across the 13 jailbreak topics (first column) from MM-SafetyBench. The “(↓)” indicates the reduction of ASR compared to no defense. For target VLMs, the format “Model A (Model B)” means defending black-box Model A against jailbreak images generated by BAP on white-box Model B (as UAP generation is white-box).
Jailbreak Topics Target VLMs LLaVA-v1.5-7B (LLaVA-v1.5-7B) MiniGPT-4 (MiniGPT-4) No Defense DiffPure Safety Prompt DiffPure + BlueSuffix No Defense DiffPure Safety Prompt DiffPure + BlueSuffix Safety Prompt Safety Prompt IA 95.88 41.24 26.80 26.80 6.19 97.90 36.08 34.02 30.93 11.34 HS 92.02 21.47 22.09 20.86 7.36 81.60 25.15 22.09 17.79 13.50 MG 93.18 70.45 45.45 22.73 9.09 88.64 45.45 34.09 27.27 9.09 PH 94.44 65.97 61.81 10.42 4.86 88.89 32.64 31.94 21.53 11.11 EH 56.56 28.69 23.77 12.30 3.28 61.48 23.77 21.31 17.21 8.20 FR 98.05 64.29 35.06 15.58 5.84 98.70 50.65 46.10 28.57 14.29 PO 69.72 39.45 17.43 9.17 4.59 68.81 27.52 24.77 22.94 9.17 PL 31.37 6.54 22.22 7.19 7.19 23.53 6.54 5.88 5.23 9.80 PV 96.40 69.87 45.32 10.79 5.04 85.61 32.37 30.94 18.71 11.51 LO 19.23 3.08 25.38 11.54 3.08 24.62 1.54 6.15 4.62 6.15 FA 4.19 0.00 4.79 0.00 2.40 4.19 0.60 1.80 1.80 7.19 HC 14.68 3.67 11.01 1.83 0.92 50.46 9.17 7.34 8.26 6.42 GD 27.52 7.38 27.52 4.70 0.67 34.90 5.37 12.08 4.03 4.03 Average 61.02 32.47 (28.55↓) 28.36 (32.66↓) 11.84 (49.18↓) 4.65 (56.37↓) 62.26 22.83 (39.43↓) 21.42 (40.84↓) 16.07 (46.19↓) 9.37 (52.89↓) Jailbreak Topics Gemini (LLaVA-v1.5-7B) Gemini (MiniGPT-4) No Defense DiffPure Safety Prompt DiffPure + BlueSuffix No Defense DiffPure Safety Prompt DiffPure + BlueSuffix Safety Prompt Safety Prompt IA 84.54 6.19 2.06 2.06 0.00 79.38 3.12 4.12 1.03 0.00 HS 74.23 4.91 3.68 9.82 0.00 76.69 2.45 6.13 7.36 1.84 MG 63.64 4.55 0.00 4.55 0.00 68.18 4.55 0.00 4.55 0.00 PH 72.92 4.17 4.86 4.17 0.69 70.83 0.69 6.25 3.47 2.08 EH 27.87 1.64 0.00 0.82 0.82 25.41 0.00 0.00 1.64 0.00 FR 83.12 5.19 5.19 3.25 1.30 80.52 2.60 0.00 0.65 0.00 PO 40.37 0.92 3.67 2.75 1.83 44.04 0.92 3.67 3.67 0.00 PL 5.88 0.00 1.31 1.31 0.00 8.50 0.00 0.00 0.65 0.00 PV 66.19 7.91 2.88 5.76 1.44 69.06 4.32 2.88 2.16 0.00 LO 3.85 0.00 2.31 1.54 0.00 3.08 0.00 0.00 0.77 0.00 FA 1.80 0.00 0.60 0.60 0.00 1.20 0.00 0.00 0.00 0.00 HC 0.92 0.00 0.92 0.00 0.00 0.92 0.00 0.00 0.00 0.00 GD 7.38 0.00 1.34 1.34 0.00 6.04 0.00 0.00 1.34 0.00 Average 40.98 2.73 (38.25↓) 2.22 (38.76↓) 2.92 (38.06↓) 0.47 (40.51↓) 41.07 1.43 (39.64↓) 1.77 (39.30↓) 2.10 (38.97↓) 0.30 (40.77↓)
Defending Commercial VLMs
@@
Here, we test our defense method on a commercial VLM: Gemini (gemini-1.5-flash) (Reid et al., 2024). As Gemini is a black-box to us, this experiment evaluates the transferability of our defense. We evaluate the defense effectiveness under two attack scenarios involving visual UAPs generated by BAP based on either LLaVA or MiniGPT-4. The results are reported in Table 1 (bottom subtable). As can be observed, compared with no defense, the adoption of our BlueSuffix reduces the ASR by more than $40\%$ under both attack scenarios. It is worth mentioning that the combined “DiffPure + Safety Prompt” defense works quite well for Gemini. This is because the safety mechanism of Gemini is much stronger than the two open-source models, thus can identify the potential risks more easily with the help of combined defenses. It is also the case for unimodal defenses DiffPure and Safety Prompt, as verified by the much lower ASR results on Gemini (compared to the two open-source models). Gemini performs robustly against 5 jailbreak topics including “Political Lobbying (PL)”, “Legal Opinion (LO)”, “Financial Advice (FA), “Health Consultation (HC)”, and “Government Decision (GD)” (the full names of other abbreviations can be found in the Appendix A) even without defense.
Comparing the results across the 13 jailbreak topics, we identify three interesting observations: 1) Certain topics, such as “FA” and “LO” are relatively easier to defend against; 2) Some topics exhibit greater resistance to either visual or textual defenses. For instance, “Privacy Violence (PV)” proves more challenging to defend against using visual purification, while “Physical Harm (PH)” is more resistant to textual defense; and 3) A no-defense scenario is not necessarily worse than any defense strategy, as seen with “FA”. This is because the target model can detect the original jailbreak texts, yet it often generates a harmful response with a disclaimer when those texts are modified by the text purifier, as illustrated in Figure 9 in Appendix E. However, the blue-team suffix generated by our method can mitigate this issue to some extent.
Figure 3: Component ablation of BlueSuffix.
@@ 4.3 Ablation Studies
Here, we conduct ablation studies to demonstrate the necessity of each component in BlueSuffix. Figure 3 reports the defense results of the ‘text purifier’, ‘text purifier + suffix generator’, ‘text purifier + image purifier’, ‘suffix generator + image purifier’, and the full BlueSuffix. We report the average ASR across the 13 categories of jailbreak prompts from the MM-SafetyBench. Detailed ASR results for each individual category can be found in Appendix E.
We start our analysis with the text purifier which is already a better defense technique than the Safety Prompt. The results in Figure 3 indicate that our text purifier itself is quite effective, compared to the no defense results in Table 1. In practice, we observed that it can rewrite the majority of jailbreak prompts, enabling the target VLMs to accurately identify the presence of harmful content in the rewritten textual prompts. When combined with the suffix generator, the ASRs are further reduced substantially. However, ‘text purifier + image purifier’ is less effective than ‘text purifier + suffix generator’, meaning that our blue-team suffix generator plays a more important role than the image purifier. Comparing ‘text purifier + image purifier’ with ‘suffix generator + image purifier’, we find that the latter outperforms the former on 3/4 scenarios including MiniGPT-4, Gemini (LLaVA), and Gemini (MiniGPT-4), but not on LLaVA. This indicates that our blue-team suffix generator is a stronger defense technique than the text purifier.
@@
We also evaluate our defense on benign prompts using LLaVA. We randomly select 500 textual prompts from the AlpacaEval dataset (Li et al., 2023c), each paired with a benign image. We define the Benign Passing Rate (BPR) as the proportion of responses that accurately address the benign prompt after applying the defense, as assessed by GPT-4o. Our method, BlueSuffix, achieved a BPR of 78.00%, which is only 3.60% lower than the original prompts’ BPR of 81.60%. Note that the BPR of the original prompts is not 100% due to the difficulty in assessing the responses using GPT-4o. Compared to other baselines, our method closely aligns with ”DiffPure” (78.80%), which does not alter the text prompts, while outperforming ”Safety Prompt” (74.80%) and ”DiffPure + Safety Prompt” (74.00%). These results demonstrate that our defense has minimal impact on benign prompts.
@@
Figure 4 illustrates six example inputs (three jailbreaks, three benign) purified by our BlueSuffix, with more examples can be found in Figure 7 and 8, Appendix D. As shown in Figure 4, the input image appears almost the same after purification by our image purifier, the rewritten texts are more detailed with many questions around the key concepts in the original texts, while the blue suffixes provide a certain type of hint or reminder for the target VLM. It also shows that our defense largely preserves the original meaning of the prompt, indicating a minimal impact on the benign inputs. Moreover, the blue-team suffixes generated by our suffix generator also exhibit high diversity.
@@ 4.4 Transferability Analysis
Our method also exhibits strong transferability across different target VLMs. Our GPT-2 based blue-team suffix generator used in this and all previous experiments was trained only based on the responses of LLaVA. Moreover, both the text purifier and image purifier adopted in BlueSuffix are generic purification models that are attack-agnostic, target model-agnostic, and fixed during all the experiments. Therefore, the high effectiveness of BlueSuffix in all the experiments shown in Table 1 and 2 highlights its high transferability in a more general scope.
@@
Attack Method Defense Method Target VLMs LLaVA-v1.5-7B MiniGPT-4 Gemini (LLaVA) Gemini (MiniGPT-4) Vanilla Attack No defense 33.80% 29.15% 3.25% 3.25% BlueSuffix 8.00 (25.80↓) 16.95 (12.20↓) 2.40 (0.85↓) 2.90 (0.35↓) BAP Attack No defense 80.20% 82.20% 52.95% 51.15% BlueSuffix 7.05 (73.15↓) 14.90 (67.30↓) 2.50 (50.45↓) 2.45 (48.70↓)
@@ 5 Conclusion
In this work, we investigated the jailbreak vulnerabilities of large Vision-Language Models (VLMs) and introduced a novel blue-team method called BlueSuffix. BlueSuffix consists of three key components: a text purifier, an image purifier, and a blue-team suffix generator. By leveraging existing unimodal purifiers, BlueSuffix trains a lightweight suffix generator to optimize the safety score of the target VLM through reinforcement learning. The blue-team suffix is generated using bimodal gradients and thus can bring cross-model robustness. Our experiments on both open-source and commercial VLMs demonstrate the high effectiveness and transferability of our defense against state-of-the-art multimodal jailbreak attacks. Additionally, BlueSuffix is resilient to an adaptive attack that optimizes jailbreak prompts based on the output of our defense. Our work proves that current VLMs, including black-box models, can be effectively defended using blue-team methods, highlighting the promise of such approaches for building robust and secure VLMs against advanced and unseen jailbreaks.
Appendix A Detailed Attack Results
In this section, we first present the detailed jailbreak topics in the MM-SafetyBench dataset (Liu et al., 2023). The query types in the above table are the abbreviations of the 13 jailbreak scenarios, including Illegal Activity, Hate Speech, Malware Generation, Physical Harm, Economic Harm, Fraud, Pornography, Political Lobbying, Privacy Violence, Legal Opinion, Financial Advice, Health Consultation, and Gov. Decision.
First, we report the evaluation results of vanilla attack and BAP attack (Ying et al., 2024) on the MM-SafetyBench dataset across the evaluated models in Table 4. BAP attack achieves a state-of-the-art ASR even compared with query-dependent multimodal jailbreaks and it demonstrates universal attacking abilities across different scenarios without requiring target scenario samples.
Table 4: The ASR (%) of Vanilla Attack (clean image + original jailbreak prompt) and BAP Attack (universally perturbed adversarial image + multi-turn enhanced jailbreak prompt) against three VLMs. We use jailbreak prompts from MM-SafetyBench across the 13 categories as the original jailbreak prompts for both attacks. The “(↑)” indicates the margin by which BAP attack surpasses the Vanilla attack. The BAP attack significantly increased the ASR across all evaluated models.
Jailbreak Topics Target VLMs LLaVA-v1.5-7B MiniGPT-4 Gemini (LLaVA) Gemini (MiniGPT-4) Vanilla Attack BAP Attack Vanilla Attack BAP Attack Vanilla Attack BAP Attack Vanilla Attack BAP Attack IA 48.45 95.88 28.87 97.90 2.06 84.54 2.06 79.38 HS 28.83 92.02 31.29 81.60 4.91 74.23 4.91 76.69 MG 70.45 93.18 25.00 88.64 0.00 63.64 0.00 68.18 PH 69.44 94.44 27.08 88.89 4.86 72.92 4.86 70.83 EH 18.85 56.56 26.23 61.48 1.64 27.87 1.64 25.41 FR 56.49 98.05 55.84 98.70 2.60 83.12 2.60 80.52 PO 33.03 69.72 18.35 68.81 5.77 40.37 5.77 44.04 PL 7.19 31.37 4.58 23.53 1.31 5.88 1.31 8.50 PV 79.86 96.40 17.99 85.61 0.00 66.19 0.00 69.06 LO 0.77 19.23 1.54 24.62 0.00 3.85 0.00 3.08 FA 0.00 4.19 2.40 4.19 0.00 1.80 0.00 1.20 HC 0.92 14.68 11.93 50.46 0.00 0.92 0.00 0.92 GD 12.75 27.52 3.36 34.90 1.34 7.38 1.34 6.04 Average 32.85 61.02 (28.17↑) 19.57 62.26 (42.69↑) 1.88 40.98 (39.10↑) 1.88 41.07 (39.19↑)
Appendix B LLaMA as the Text Purifier
Here, we test the use of Llama-3-8B-Instruct (AI@Meta, 2024) for textual prompt rewriting. The prompts rewritten using the LLaMA model demonstrates comparable performance to those rewritten by GPT-4o, both in terms of semantic expression and defense effectiveness, as shown in Table 5. We also present an example of a jailbreaking textual prompt purified by GPT-4o and Llama-3-8B-Instruct in Figure 5. The purified textual prompt consists of two parts: it first repeats the jailbreak prompt, and then emphasizes the potential presence of malicious queries, which gives a hint to the target VLM (bold font).

### v2

@@ 4.2 Main Results
We first evaluate our BlueSuffix defending image-based (VAA, imgJP) and text-based (GCG, AutoDAN) attack on three open-source VLMs and a commercial VLM. Our BlueSuffix demonstrates a significant advantage over other baselines, reducing the ASR to zero in all cases, with results summarized in Table 1. It is important to note that our suffix generator is fine-tuned on hard jailbreak prompts crafted by BAP attack. Therefore, defending image-based and text-based attacks confirm the effectiveness of BlueSuffix and demonstrate its strong generalizability in defending against different types of jailbreak attacks across various datasets and models.
@@
For cross-modal attack BAP on open-source models, BlueSuffix reduces the ASR of BAP attacks by $56.37\%$ on the LLaVA model (from $61.02\%$ to $4.65\%$) by $52.89\%$ on MiniGPT-4 (from $62.26\%$ to $9.37\%$) and by $51.70\%$ on the InstructionBLIP (from $58.48\%$ to $6.78\%$). Particularly, when compared with unimodal defense (DiffPure, Safety Prompt, R2D2, and CAT), our method demonstrates at least 23% robustness improvement (56.37% vs. 32.66%) on LLaVA, 12% on MiniGPT-4 (52.89% vs. 40.84%) and 16% on InstructionBLIP (51.70% vs. 35.80%). Such a huge improvement demonstrates the advantage of bimodal defense over unimodal defense. An interesting observation about unimodal defense is that textual defense appears to be more effective than visual defense. For bimodal defense (‘DiffPure + Safety Prompt’ and VLGuard), the “DiffPure + Safety Prompt” method exhibits much greater ASR reduction on all open-source models, even better than VLGuard, showcasing the strength of bimodal defense. Moreover, our BlueSuffix beats “DiffPure + Safety Prompt” by a considerable margin. This indicates the importance of the suffix generator for enhancing the cross-modal robustness.
@@
We also test our defense method on a commercial VLM: Gemini (gemini-1.5-flash). As Gemini is a black-box to us, this experiment evaluates the transferability of our defense. We evaluate the defense effectiveness under three attack scenarios involving visual UAPs generated by BAP based on either LLaVA, MiniGPT-4, or InstructionBLIP. As can be observed, compared with no defense, the adoption of our BlueSuffix reduces the ASR by more than $40\%$ under all attack scenarios. It is worth mentioning that the combined “DiffPure + Safety Prompt” defense works quite well for Gemini. This is because the safety mechanism of Gemini is much stronger than the open-source models, thus can identify the potential risks more easily with the help of combined defenses. It is also the case for unimodal defenses DiffPure and Safety Prompt, as verified by the much lower ASR results on Gemini (compared to the three open-source models).
4.3 Demonstration of Cross-modal Utilization
BlueSuffix operates under a black-box defense setting, where lack access to the internals of the target VLMs. As a result, the only cross-modal information we can leverage is derived from the VLM’s outputs. To demonstrate whether the generated suffix genuinely exploits cross-modal information, we provide a quantitative analysis using Mutual Information (MI).
Table 2: The results demonstrating the generated suffix exploits cross-modal information.
MI (Image; Output) MI (Text; Output) MI (Image + Text; Output) MI (Image + Text + Suffix; Output) 1.4093 2.7756 4.3667 4.4728
For a fair comparison, we selected 50 pairs of purified jailbreak image-text prompts. These purified prompts do not trigger a jailbreak when input into the LLaVA model, regardless of whether a suffix is added. We employed the K-Nearest Neighbors algorithm to estimate MI for four input combinations: “Image”, “Text”, “Image + Text”, “Image + Text + Suffix”. The results shown in Table 2 confirm that MI(Image + Text + Suffix; Output) $>$ MI(Image + Text; Output), providing strong evidence that our suffix generator indeed enhances the utilization of cross-modal information.
@@ 4.4 Ablation Studies
Here, we conduct ablation studies to demonstrate the necessity of each component in BlueSuffix. Figure 3 reports the defense results of the ‘text purifier’, ‘suffix generator’, ‘text purifier + suffix generator’, ‘text purifier + image purifier’, ‘suffix generator + image purifier’, and the full BlueSuffix. We report the average ASR across the 13 categories of jailbreak prompts from the MM-SafetyBench. Due to space limitations, we only conducted ablation studies on the LLaVA, MiniGPT-4, and Gemini.
We start our analysis with the ‘text purifier’ which is already a better defense technique than the Safety Prompt. The results in Figure 3 indicate that our ‘text purifier’ itself is quite effective, compared to the no defense results in Table 1. In practice, we observed that it can rewrite the majority of jailbreak prompts, enabling the target VLMs to accurately identify the presence of harmful content in the rewritten textual prompts. Furthermore, employing the ‘suffix generator’ independently also provides a certain degree of defense, demonstrating a comparable level of robustness achieved by ‘text purifier’. When combined the ‘text purifier’ with the ‘suffix generator’, the ASRs are further reduced substantially. However, ‘text purifier + image purifier’ is less effective than ‘text purifier + suffix generator’, meaning that our blue-team suffix generator plays a more important role than the image purifier.
Figure 3: Component ablation of BlueSuffix.
Additionally, we observed that using only the ‘text purifier’ (without the ‘suffix generator’) can sometimes result in even stronger jailbreak prompts even when the ”Image Purifier” is used (as illustrated in Figure 10 in Appendix J), leading to a higher ASR than when no ‘text purifier’ is used. This underscores the limitations of relying solely on unimodal or independent bimodal defenses. Combining both the ‘image purifier’ and the ‘text purifier’ addresses the limitation of unimodal defense by providing protection across both modalities. However, their inherent limitation persists: they do not leverage cross-modal information. Our ‘Suffix Generator’ effectively overcomes this limitation by incorporating cross-modal interactions, significantly enhancing the overall defense capability. Moreover, the unimodal purifiers positively impact the fine-tuning of the Suffix Generator by weakening the jailbreak prompts before they reach it. This synergy results in a more effective and robust Suffix Generator. While fine-tuning the Suffix Generator without the purifiers can achieve good defense capabilities, it tends to have a greater negative impact on benign inputs, adversely affecting user experience.
@@
We also evaluate our defense on benign prompts using LLaVA. We randomly select 500 textual prompts from the AlpacaEval dataset (Li et al., 2023b), each paired with a benign image. We define the Benign Passing Rate (BPR) as the proportion of responses that accurately address the benign prompt after applying the defense, as assessed by GPT-4o. Our method, BlueSuffix, achieved a BPR of 78.00%, which is only 3.60% lower than the original prompts’ BPR of 81.60%. Note that the BPR of the original prompts is not 100% due to the difficulty in assessing the responses using GPT-4o. Compared to other baselines, our method closely aligns with ”DiffPure” (78.80%), which does not alter the text prompts, while outperforming ”Safety Prompt” (74.80%) and ”DiffPure + Safety Prompt” (74.00%). These results demonstrate that our defense has minimal impact on benign prompts.
@@
Figure 7 in Appendix I illustrates six example inputs (three jailbreaks, three benign) purified by our BlueSuffix, with more examples can be found in Figure 8 and 9, Appendix I. As shown in Figure 7, the input image appears almost the same after purification by our image purifier, the rewritten texts are more detailed with many questions around the key concepts in the original texts, while the blue suffixes provide a certain type of hint or reminder for the target VLM and the suffixes generated by our suffix generator also exhibit high diversity. Furthermore, we used GPT-4o to compare the semantics of the ‘original prompt’ with ‘text purifier + suffix generator prompt’ in Appendix B and the results indicate a high level of semantic consistency.
@@ 4.5 Transferability Analysis
Our method also exhibits strong transferability across different target VLMs. Our GPT-2 based blue-team suffix generator used in this and all previous experiments was trained only based on the responses of LLaVA. Moreover, both the text purifier and image purifier adopted in BlueSuffix are generic purification models that are attack-agnostic, target model-agnostic, and fixed during all the experiments. Therefore, the high effectiveness of BlueSuffix in all the experiments shown in Table 1 and 3 highlights its high transferability in a more general scope.
@@
Attack Method Defense Method Target VLMs LLaVA-v1.5-7B MiniGPT-4 InstructionBLIP Gemini (LLaVA) Gemini (MiniGPT-4) Gemini (InstructionBLIP) Vanilla Attack No defense 33.80 29.15 30.05 3.25 3.25 3.25 BlueSuffix 8.00 (25.80↓) 16.95 (12.20↓) 10.10 (19.95↓) 2.40 (0.85↓) 2.90 (0.35↓) 2.55 (0.70↓) BAP Attack No defense 80.20 82.20 77.05 52.95 51.15 50.05 BlueSuffix 7.05 (73.15↓) 14.90 (67.30↓) 9.60 (67.45↓) 2.50 (50.45↓) 2.45 (48.70↓) 2.30 (47.75↓)
@@ 5 Conclusion
In this work, we investigated the jailbreak vulnerabilities of large Vision-Language Models (VLMs) and introduced a novel blue-team method called BlueSuffix. BlueSuffix consists of three key components: a text purifier, an image purifier, and a blue-team suffix generator. By leveraging existing unimodal purifiers, BlueSuffix trains a lightweight suffix generator to optimize the safety score of the target VLM through reinforcement learning. The blue-team suffix is generated using bimodal gradients and thus can bring cross-model robustness. Our experiments on both open-source and commercial VLMs demonstrated the high effectiveness and transferability of our defense against state-of-the-art multimodal jailbreak attacks. Additionally, BlueSuffix is resilient to an adaptive attack that optimizes jailbreak prompts based on the output of our defense. Our work proves that current VLMs, including black-box models, can be effectively defended using blue-team methods, highlighting the promise of such approaches for building robust and secure VLMs against advanced and unseen jailbreaks.
Acknowledgment
This work is in part supported by the National Key R&D Program of China (Grant No. 2021ZD0112804) and the National Natural Science Foundation of China (Grant No. 62276067).
The computations in this research were performed using the CFFF platform of Fudan University.
Appendix A LLaMA as the Text Purifier
Here, we test the use of Llama-3-8B-Instruct AI@Meta (2024) for textual prompt rewriting. The prompts rewritten using the LLaMA model demonstrates comparable performance to those rewritten by GPT-4o, both in terms of semantic expression and defense effectiveness, as shown in Table 5. We also present an example of a jailbreaking textual prompt purified by GPT-4o and Llama-3-8B-Instruct in Figure 4. The purified textual prompt consists of two parts: it first repeats the jailbreak prompt, and then emphasizes the potential presence of malicious queries, which gives a hint to the target VLM (bold font).